# Structured and Balanced Multi-component and Multi-layer Neural Networks

### v2

@@ 6.1 Advantages compared to FCNNs or MLPs
The two key differences between a standard FCNN or MLP and an MMNN are 1) the introduction of the weights ${\bm{A}},{\bm{c}}$ for different linear combinations of hidden neurons (or perceptrons) as the multi-components in each layer, and 2) the training strategy that fixes those randomly initialized ${\bm{W}},{\bm{b}}$ (random features) in the hidden neurons. Hence it is extremely easy to modify a FCNN or MLP to an MMNN.
MMNNs are much more effective than FCNNs in terms of representation, training, and accuracy especially for complex functions. In comparison, as shown in those experiments in Section 3, MMNNs 1) have much fewer training parameters, 2) converge much faster in training, 3) achieve much better accuracy. Moreover, experiments show that training process of MMNNs converges not only faster but also with a steady rate while FCNNs saturates pretty early to a quite low accuracy, as commonly observed in practices. These nice behaviors of MMNNs are due to their balanced structure for smooth decomposition as well as the training strategy. In practice, the introduction of ${\bm{A}},{\bm{c}}$ in MMNNs provides an important balance between the network width, which is the number of hidden neurons (basis functions) and can be very large, and the dimension of the input space, which is the number of components from the previous layer and can be much smaller than the network width. In other words, using a few linear combinations of the basis functions can capture smooth structures in the input space well. On the other hand, for FCNNs the two are the same and no balance is exerted.
6.2 Practical guidelines for MMNN
@@
1. The network width should provide enough resolution to capture fine details of the target function. This means that the width should be at least comparable to the size of an adaptive mesh that can approximate the target function well.
2. The number of components (rank) is related to the overall complexity of the target function which depends on its spatial domain and Fourier domain representation as well as the input dimension. As indicated by our mathematical multi-component construction, it is related to the “divide and conquer” strategy.
3. The number of layers (depth) is also related to the overall complexity of the target function as for the number of components. Rank and depth are complementary but work together effectively for a smooth decomposition of the target function. The rule of thumb for depth is similar to that for the rank.
@@ Acknowledgments
S. Zhang was partially supported by the start-up fund P0053092 from The Hong Kong Polytechnic University. H. Zhao was partially supported by NSF grants DMS-2309551, and DMS-2012860. Y. Zhong was partially supported by NSF grant DMS-2309530, H. Zhou was partially supported by NSF grant DMS-2307465.
Appendix A Tests in three dimension and higher
In this section, we test a few examples in three and four dimensions. Even sampling an interesting function becomes challenging as the dimension becomes higher. Although our examples are limited by our computation power using a laptop, our tests show that MMNN performs well and is more effective than a fully connected network.
The first example is a 3D function a level set of which is shown in Figure 35. Using polar coordinates $(r,\theta,\phi)$, $\theta\in[0,\pi]$, $\phi\in[0,2\pi)$, the target function $f(x,y,z)$ is defined as:
where
Figure 33: Surface of the levelset $f(r,\theta,\phi)=0.5$.
Figure 34: Surface of the levelset $h(r,\theta,\phi)=0.5$.
Figure 35: Training and test errors (MSE) vs. epoch.
Our MMNN is of a compact size $(600,20,8)$. For this test, a total of $111^{3}$ data are sampled on a uniform grid in $[-1,1]^{3}$ with a mini-batch size of $999$ and a learning rate of $0.0005\times 0.9^{\lfloor k/6\rfloor}$ for epochs $k=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},300$. Figure 35 gives the error plot. As shown in Figures 35 and 35, the levelsets corresponding to the target function $f$ and the learned MMNN approximation $h$ are nearly identical. To visually demonstrate the quality of the approximation and complex structure of the 3D function, we present several slices of the target function and the MMNN approximation by fixing either $x$, $y$, or $z$ in Figure 36.
(a) $z=0$.
(b) $z=0.1$.
(c) $y=0.2$.
(d) $y=0.3$.
(e) $x=0.4$.
(f) $x=0.5$.
Next, we consider the probability density function (PDF) of a Gaussian (normal) distribution in 4D,
where ${\bm{\Sigma}}$ is the covariance matrix. We set $\bm{\mu}=\bm{0}$ and
Figure 38: Training and test errors (MSE) vs. epoch.

### v3

@@ 6.1 Advantages compared to FCNNs or MLPs
The two key differences between a standard FCNN or MLP and an MMNN are (1) the introduction of the weights ${\bm{A}},{\bm{c}}$ for different linear combinations of hidden neurons (or perceptrons) as the multi-components in each layer, and (2) the training strategy that fixes those randomly initialized ${\bm{W}},{\bm{b}}$ (random features) in the hidden neurons. Hence it is extremely easy to modify a FCNN or MLP to an MMNN.
MMNNs are much more effective than FCNNs in terms of representation, training, and accuracy especially for complex functions. In comparison, as shown in those experiments in Section 3, MMNNs (1) have much fewer training parameters, (2) converge much faster in training, (3) achieve much better accuracy. Moreover, experiments show that training process of MMNNs converges not only faster but also with a steady rate while FCNNs saturates pretty early to a quite low accuracy, as commonly observed in practices. These nice behaviors of MMNNs are due to their balanced structure for smooth decomposition as well as the training strategy. In practice, the introduction of ${\bm{A}},{\bm{c}}$ in MMNNs provides an important balance between the network width, which is the number of hidden neurons (basis functions) and can be very large, and the dimension of the input space, which is the number of components from the previous layer and can be much smaller than the network width. In other words, using a few linear combinations of the basis functions can capture smooth structures in the input space well. On the other hand, for FCNNs the two are the same and no balance is exerted.
6.2 Practical guidelines for MMNNs
@@
  1. 1.
    The network width should provide enough resolution to capture fine details of the target function. This means that the width should be at least comparable to the size of an adaptive mesh that can approximate the target function well.
  2. 2.
    The number of components (rank) is related to the overall complexity of the target function which depends on its spatial domain and Fourier domain representation as well as the input dimension. As indicated by our mathematical multi-component construction, it is related to the “divide and conquer” strategy.
  3. 3.
    The number of layers (depth) is also related to the overall complexity of the target function as for the number of components. Rank and depth are complementary but work together effectively for a smooth decomposition of the target function. The rule of thumb for depth is similar to that for the rank.
@@ Acknowledgments
S. Zhang was partially supported by the start-up fund P0053092 from The Hong Kong Polytechnic University. H. Zhao was partially supported by NSF grants DMS-2309551, and DMS-2012860. Y. Zhong was partially supported by NSF grant DMS-2309530. H. Zhou was partially supported by NSF grant DMS-2307465.